(ax^2)-bx+c=0 has distinct roots lyin b/w 0 & 1.then which of the following holds if a,b,c E N.



  1. abc >=20

  2. abc >=25

  3. abc <=16

  4. abc =9.

Grade:11

1 Answers

basit ali
12 years ago

Here, b/a=0+1=1, c/a=0*1≠0 ( here 0 is not exactly 0) 

 b=a, c=na(n is any integer )

 let, ax2-ax+na=0

  a(x2-x+n)=0

  x2-x+n=0

 x2=x

 12=1. so x=1 is true. 

 a-b+c=0

 -bx+na= 0

-bx=-na

 b=na

  hit and trial, b=a=√3 and c=3 .then,

abc = 9. this is the answer.  plz aprove the answer.